OpenStudy (anonymous):

5x^-2y^10 over 2x^-1(-3x^-3y^-1)^-2

OpenStudy (turingtest):

\[{5x^{-2}y^{10}\over2x^{-1}(-3x^{-3}y^{-1})^{-2}}\]is the problem?

OpenStudy (anonymous):

Yes

OpenStudy (turingtest):

three rules to use here:\[(x^a)^b=x^{ab}\]\[x^a\cdot x^b=x^{a+b}\]\[{x^a\over x^b}=x^{a-b}\]use these rules on all like terms carefully and you will get the answer

OpenStudy (anonymous):

Ok Thanks And Sorry About The other Post..

OpenStudy (turingtest):

No problem, you're new again, welcome to OS, let me know if you need more help

OpenStudy (anonymous):

Ok:)